# SIToBI - A Speech Prosody Annotation Tool for Indian Languages

### v1

@@ III-B Accuracy of Segmentation
The accuracy with which the tool segments the given speech signals using the HMM-based forced-Viterbi alignment procedure is evaluated at the phoneme level, as syllable and word-level segmentations are derived from the phoneme-level segmentation. To evaluate the accuracy, the segmentation error is calculated as the time difference between the manually derived duration for each phoneme and the corresponding duration obtained from the tool. The distribution of the segmentation errors for each language when language-dependent and language-independent models are used is shown in Fig. 4.
@@
It can be observed that ninety percent of segments in the language-dependent models across all three languages exhibit errors of less than 10ms . On average, the error lies between 10 to 30ms for the three languages when using both language-dependent and language-independent models. There is marginal difference of 10% error observed in language-independent models compared to language-dependent models arises from the use of shared phonetic models across languages, which leads to some generalization errors. Nevertheless, this trade-off is deemed acceptable given the increased flexibility and efficiency offered by language-independent models, which are essential for accommodating multiple languages effectively.
(a) Tamil: Independent and Dependent
(b) Hindi: Independent and Dependent
(c) English:Independent and Dependent
@@
Figure 4: Comparison of Segmentation Error for Language-Independent and Language-Dependent Models (Tamil, Hindi, English)
III-C Accuracy of Break Indices
@@
To assess the accuracy of break indices derived from the tool, these were compared with the manual annotations. As shown in Table I, the break indices are derived with an accuracy of around 95%. It is observed that in the remaining 5% of the silence regions, the boundaries detected based on the spectral flatness is inaccurate resulting mostly in the incorrect identification of a break index of 2 as 3 or 1. Further, some of the silence regions occur before or between stops consonants and fricatives were not considered as silence. This is reflected in the confusion matrices in Fig. 5. The lower accuracy in Hindi compared to Tamil and English can be attributed to the influence of fricatives. Silence preceding fricatives tends to have a lower spectral flatness (SF) value, which may contribute to the decreased accuracy in Hindi.
TABLE I: Overall Accuracy of Break Indices for three Languages
Language Overall Accuracy (%)
Hindi 91.53
English 97.62
Tamil 98.56
@@
Figure 5: Confusion Matrices for the Identification of Break Indices: (a) Tamil, (b) English, (c) Hindi
III-D Accuracy of Pitch Contour Labeling
@@
As in the previous sections, the accuracy of the pitch contour labels is assessed by comparing the manual annotations with those derived from the tool. Manual labeling involves a careful analysis of the audio signals, where expert annotators listen to the speech and identify pitch contours by visually examining the waveform and their pitch frequency. The annotators mark significant points, such as rise and fall, to capture the contour patterns present in the speech. It is observed that the designed rules capture the shape of the pitch contours with an accuracy of approximately 99%, as summarized in Table II. The LLH label has a high error rate for English, while the L label has a high error rate for both Tamil and Hindi.
TABLE II: Accuracy of Pitch Labels for Three Languages
Language Accuracy of pitch labels (%)
Tamil 99.91
Hindi 99.78
English 97.24
@@
IV Influence of pitch contour on Language Identification Using the SIToBI Tool
@@
This study aims to check whether the frequency of pitch contours has any influence on language identification. Specifically, the current work attempts to identify the language of a given word based on syllable-level pitch contours, as prosodic information varies across languages. For this task, Tamil and Hindi are considered. Initially, monosyllabic words from the corpus described in Section III-A are used, comprising 1500 monosyllabic words from both Tamil and Hindi data. Pitch contour labels are estimated from 90% of these words, and the normalized frequency of occurrence for each contour in each language is computed. For the remaining 10% of the words, the previously computed normalized frequency of occurrence is assigned as a score based on the pitch contour.
@@
Figure 6 illustrates the accuracy of language identification for given words. It is observed that identifying Hindi words becomes easier as the number of syllables increases. Only around 52% of monosyllabic words are identified correctly, while 73% to 82% of trisyllabic, quadsyllabic, and pentasyllabic words are accurately identified. In contrast, Tamil shows slightly lower identification accuracy for trisyllabic and quadsyllabic words, around 65%, whereas other word lengths achieve an accuracy exceeding 72%.
@@
V Conclusion
The proposed SIToBI tool provides rich prosodic features for three languages, namely Indian English, Tamil, and Hindi. The performance of each module of the tool has been compared with manual annotations, and the accuracy has been computed, demonstrating its effectiveness in capturing phonetic and prosodic details. While the current work focused on these three languages, the tool can be easily extended to other languages by retraining the phoneme HMMs and incorporating automatic speech recognition (ASR) systems for the new languages. Additionally, the modular design of the tool allows for adaptability to various prosodic feature extraction tasks, making it a versatile tool for speech processing across different linguistic contexts. The general syllabification algorithm used in both language-dependent and language-independent models ensures that the tool can handle diverse language structures with minimal degradation in performance. Future work could explore applying this tool to low-resource languages, as well as refining the pitch contour labeling for better accuracy in tonal languages. The SIToBI tool has the potential to significantly contribute to multilingual speech processing and analysis in a variety of applications, from language identification to speech synthesis and recognition.

### v2

@@ 3.2 Accuracy of Segmentation
The accuracy with which the tool segments the given speech signals using the HMM-based forced-Viterbi alignment procedure is evaluated at the phoneme level, as syllable and word-level segmentations are derived from the phoneme-level segmentation. To evaluate accuracy, the segmentation error is calculated as the time difference between the manually derived duration for each phoneme and the corresponding duration obtained from the tool. The distribution of the segmentation errors for each language when language-dependent and language-independent models are used is shown in figure 4.
@@
It can be observed that ninety percent of segments in the language-dependent models across all five languages exhibit errors of less than 10ms. On average, the error lies between 10 to 30ms for the five languages when using both language-dependent and language-independent models. There is marginal difference of 10% error observed in language-independent models compared to language-dependent models arises from the use of shared phonetic models across languages, which leads to some generalization errors. Nevertheless, this trade-off is deemed acceptable given the increased flexibility and efficiency offered by language-independent models, which are essential for accommodating multiple languages effectively.
@@
Figure 4: Comparison of Segmentation Error for Language-Independent and Language-Dependent Models
3.3 Accuracy of Break Indices
@@
To assess the accuracy of break indices derived from the tool, these were compared with the manual annotations. As shown in Table 1, the break indices are derived with an accuracy of around 95%. It is observed that in the remaining 5% of the silence regions, the boundaries detected based on the spectral flatness is inaccurate resulting mostly in the incorrect identification of a break index of 2 as 3 or 1. Further, some of the silence regions occur before or between stops consonants and fricatives were not considered as silence. This is reflected in the confusion matrices in figure 5. While Tamil, Indian English, Gujarati, and Malayalam exhibit comparable performance, slightly lower accuracy is observed for Hindi attributed to the influence of fricatives. Silence preceding fricatives tends to have a lower Spectral Flatness (SF) value, which may contribute to the decreased accuracy in Hindi.
Refer to caption      Refer to caption
@@
Figure 5: Confusion matrices for the identification of Break Indices: (a) Tamil, (b) Indian English, (c) Hindi, (d) Gujarati, and (e) Malayalam.
3.4 Accuracy of Pitch Contour Labeling
@@
As in the previous sections, the accuracy of the pitch contour labels is assessed by comparing the manual annotations with those derived from the tool. Manual labeling involves a careful analysis of the audio signals, where expert annotators listen to the speech and identify pitch contours by visually examining the waveform and their pitch frequency. The annotators mark significant points, such as rise and fall, to capture the contour patterns present in the speech. It is observed that the designed rules capture the shape of the pitch contours with an accuracy of approximately 99%, as summarized in Table 1. The LLH label has a high error rate for English, while the L label has a high error rate for both Tamil and Hindi. Gujarati shows relatively higher confusion for the HL label, whereas Malayalam exhibits higher errors for the LH label.
3.5 Performance Across Languages and Noise Conditions
To analyze the impact of different training data sources, we consider three individual training conditions: TTS-only, ASR-only, and noise-augmented-only training. Models trained exclusively on clean TTS data achieve slightly higher clean-speech performance, with language-dependent phoneme segmentation accuracies of approximately 95–96%, pitch contour correlations close to 0.99, and break index F1-scores in the range of 0.96–0.97, but with reduced robustness under noisy conditions. ASR-only training yields moderate and stable performance (Seg-LD $\approx$ 92–93%, BI-F1 $\approx$ 0.93–0.94), while noise-augmented-only training improves robustness at the cost of a small reduction in clean prosodic accuracy (Seg-LD $\approx$ 90–92%, BI-F1 $\approx$ 0.91–0.93). The combined training strategy adopted in this work provides a balanced trade-off between clean-speech accuracy and noise robustness, resulting in the stable performance reported in Table 1. As expected, noise-augmented training results in a minor degradation in clean-speech accuracy, offset by improved robustness under noisy conditions. The results reported in Table 1 correspond to models trained using a combined dataset consisting of 1 hour of clean TTS speech, 1 hour of multispeaker ASR speech, and 2 hours of noise-augmented speech, as described in 3.1.
Abbreviations: W/Utt = average words per utterance; Seg-LD = language-dependent phoneme segmentation accuracy (%); Seg-LI = language-independent phoneme segmentation accuracy (%); Pitch-$r$ = pitch contour correlation; BI-F1 = Break Index F1-score.
Lang. Hours #Utts W/Utt Seg-LD Seg-LI Pitch-$r$ BI-F1
Eng 4T/2E 4200 6.4 94.6 85.1 0.98 0.95
Tam 4T/2E 4100 5.9 95.2 86.0 0.98 0.96
Hin 4T/2E 4050 6.7 91.8 82.4 0.97 0.91
Guj 4T/2E 3950 6.1 93.7 84.6 0.98 0.96
Mal 4T/2E 4000 5.6 94.1 85.0 0.98 0.93
Table 1: Prosody accuracy across languages with noise-augmented training data.
@@ 3.5 Performance Across Languages and Noise Conditions
4 Influence of pitch contour on Language Identification Using the AutoProsody Tool
@@
This study aims to check whether the frequency of pitch contours has any influence on language identification. Specifically, the current work attempts to identify the language of a given word based on syllable-level pitch contours, as prosodic information varies across languages. For this task, Tamil, Hindi, Indian English are considered. Initially, monosyllabic words from IITM IndicTTS corpus[8] , comprising 1500 monosyllabic words from the Tamil, Hindi, Indian English data. Pitch contour labels are estimated from 90% of these words, and the normalized frequency of occurrence for each contour in each language is computed. For the remaining 10% of the words, the previously computed normalized frequency of occurrence is assigned as a score based on the pitch contour.
@@
Figure 6 illustrates the precision of language identification in different syllable lengths for Hindi, Tamil, and Indian English. The results reveal that language identification accuracy generally improves with an increase in the syllable count. For Hindi, the model struggles with monosyllabic and disyllabic words, achieving only 13.3% and 4.5% accuracy, respectively. However, performance improves significantly for longer words, with accuracy reaching 41.1% for trisyllabic, 66.7% for quadsyllabic, and 76.7% for pentasyllabic words. Tamil shows a more balanced trend, maintaining an accuracy greater than 23% even for monosyllabic words and steadily increasing to 79. 0% for quadsyllabic words and 75. 0% for pentasyllabic words. English shows a distinct pattern, with high accuracy for disyllabic words at 76.0%, moderate accuracy for tri- and pentasyllabic words (55.9% and 72.4% respectively), and relatively low performance for monosyllabic (18.8%) and quadsyllabic (54.3%) words. These patterns suggest that syllable length plays a crucial role in the model’s ability to correctly identify the language, likely because of the increased linguistic cues available in longer words.
@@
5 Conclusion
The proposed AutoProsody tool provides rich prosodic features for five languages, namely Indian English, Tamil, Hindi, Malayalam and Gujarati. The performance of each module of the tool has been compared with manual annotations, and the accuracy has been computed, demonstrating its effectiveness in capturing phonetic and prosodic details. While the current work focused on these five languages, the tool can be easily extended to other languages by retraining the phoneme HMMs and incorporating automatic speech recognition (ASR) systems for the new languages. Additionally, the modular design of the tool allows for adaptability to various prosodic feature extraction tasks, making it a versatile tool for speech processing across different linguistic contexts. The general syllabification algorithm used in both language-dependent and language-independent models ensures that the tool can handle diverse language structures with minimal degradation in performance. Future work could explore applying this tool to low-resource languages, as well as refining the pitch contour labeling for better accuracy in tonal languages. The AutoProsody tool has the potential to significantly contribute to multilingual speech processing and analysis in a variety of applications, from language identification to speech synthesis and recognition.
Acknowledgments
The current work is carried out as a part of the project titled, “Prosody Modeling”, under the sub-project of the NLTM BHASHINI project, titled, “Speech technologies in Indian languages”, funded by the Ministry of Electronics and Information Technology, Government of India, with reference number, 11(1)/2022-HCC(TDIL).